Derive the cost function for AT&T/Sprint post-merger for any value of x. suppose that service quality declines by one-third to s = 8 following the merger. In addition, assume that the market for long-distance telephone service remains perfectly competitive following the merger. Determine the values of x for which the DOJ will approve this merger
1
Expert's answer
2020-05-27T10:00:19-0400
"Q=x(K+\frac{1}{2}L)"


"Q=a-bp+s"

s=8



"a-bp+8=x(K+\frac{1}{2}L)"


"x=\frac{a-bp+8}{K+\frac{1}{2}L}"